OpenStudy (anonymous):

find the two unknown lengths for: A right triangle has a hypotenuse that is 3 feet longer than one leg. The other leg is 4 feet.

OpenStudy (pfenn1):

Well, you know that for a right triangle\[a ^{2}+b ^{2}=c ^{2}\] where a and b are the lengths of the sides and c is the length of the hypotenuse. If x is the unknown length of one side, how would you express the length of the hypotenuse?

OpenStudy (anonymous):

\[x+3=\sqrt{4^{2}+x ^{2}}\]

OpenStudy (anonymous):

3x=8+x^2?

OpenStudy (pfenn1):

No. The length of the hypotenuse would be x+3. so you would have \[4^{2}+x ^{2}=(x+3)^{2}\]\[4^{2}+x ^{2}=x ^{2}+6x=9\]Can you solve this equation for x?

OpenStudy (anonymous):

that would be the one...ehehe

OpenStudy (pfenn1):

that should have been \[4^{2}+x ^{2}=x ^{2}+6x+9\]

OpenStudy (anonymous):

x=7/6

OpenStudy (pfenn1):

yep, that is what I got.

OpenStudy (anonymous):

yahp

OpenStudy (anonymous):

That's it?

OpenStudy (anonymous):

yes...simple

OpenStudy (pfenn1):

well, that gave you the length of the side, x remember that the hypotenuse is x+3

OpenStudy (anonymous):

then find the lenght of your hypotenuse which is x+3

OpenStudy (anonymous):

25/6?

OpenStudy (pfenn1):

yes

OpenStudy (anonymous):

so 7/6, 25/6?

OpenStudy (pfenn1):

That is the answer I got,

OpenStudy (anonymous):

Yaaaaay! Thanks @pfenn1 @ren

OpenStudy (pfenn1):

you're welcome

OpenStudy (anonymous):

welcome